The electrical resistance of a platinum wire of zero degree celsius is 9.8 ohmes and 11.0 ohmes at 100 degree celsius. calculate the temperature when the resistance is 10.4 ohms assuming that the change in electrical resistance of the conductor is proportional to the change in temperature


1
Expert's answer
2022-01-23T16:13:13-0500

Solution

The temperature dependency of resistance is given by

"R_t=R_0(1+\alpha\Delta T)"

0 means at zero celsius temperature

t means at Any other temperture

So

Constant

"\alpha =\frac{R_t-R_0}{\Delta T}"


Putting all values

"\alpha =\frac{11-9.8}{100}=0.012 \Omega/°C"

Now this temperature is T

Then resistance is 10.4 "\Omega"

"T=\frac{R_t-R_0}{\alpha}=\frac{10.4-9.8}{0.012}=50°C"